OpenStudy (anonymous):

What are the approximate solutions of 2x^2 + 9x = 8 to the nearest hundredth?

jhonyy9 (jhonyy9):

x_1,2=(-9 +/- sqrt(81+64)/4

jhonyy9 (jhonyy9):

- can you continue it ?

OpenStudy (anonymous):

i'm stuck there

OpenStudy (anonymous):

Ok I got the answer thanks for everyone's help!

jhonyy9 (jhonyy9):

yw good luck bye

OpenStudy (anonymous):

is it the quadratic equation? ... it seems to me like it is 2 *x^2+9x-8=0 .. ?

OpenStudy (anonymous):

\[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] ... ..